plant to absorb the backlog of demand that now exists. The primary location being considered will have fixed costs of $7,400 per month and variable costs of $0.73 per unit produced. Each item is sold to retailers at a price that averages $1.13 a) The volume per month is required in order to break even = b) The profit or loss would be realized on a monthly volume of 61,000 units = c) The volume is needed to obtain a profit of $16,000 per month d) The volume is needed to provide revenue of $23,000 per month = (in whole number) (in whole (in whole number)
Answer #1

Answer

  • All working forms part of the answer itself.
  • Amounts are in “$”
  • Working

A

Selling price per unit

$1.13

B

Variable cost per unit

$0.73

C=A-B

Contribution margin per unit

$0.4

D

Fixed Cost

$7400

  • Answer a

A

Fixed Cost

$7400

B

Contribution margin per unit

$0.4

C=A/B

Break Even in Volume (units)

18500

  • Answer b

A

Contribution margin

[61000 x $0.4] $24400

B

Fixed Cost

$7400

C=A-B

Profit realized on sale of 61000 units

$17000

  • Answer c

A

Required profit

$16000

B

Fixed Cost

$7400

C=A+B

Contribution margin required

$23400

D

Contribution margin per unit

$0.4

E=C/D

Volume (units) required to earn $16000 profit

58500

  • Answer d

A

Sales revenue

$23000

B

Selling price per unit

$1.13

C=A/B

Volume required to provide revenue of $23000

20354